How much water at (20°C) is required to cool a 5kg burning wood log from 300°C to 80°C ?


1
Expert's answer
2021-02-01T15:07:31-0500
"t_1=80^0"


"t_2=300^0"


"t_3=20^0"

"c_cm_c(t_2-t_1)=\lambda m_w+c_wm_w(t_1-t_2)"

"m_w=\frac {1.7\times 10^3\times 5 \times 220^0}{2.3\times 10^6+4.2\times 10^6\times 60^0}=0.732 kg"